OpenStudy (anonymous):

What is the equation of the line, in point-slope form, that passes through the points (9, 1) and (4, 16)

OpenStudy (anonymous):

What is the equation of the line, in point-slope form, that passes through the points (9, 1) and (4, 16)

OpenStudy (triciaal):

first find the slope slope,m = change in y / corresponding change in x

Nnesha (nnesha):

y=-3x+28

OpenStudy (triciaal):

what do you get for the slope?

OpenStudy (triciaal):

@Nnesha the question said in point-slope form,

OpenStudy (triciaal):

that's not what I got, i have to check mine

OpenStudy (triciaal):

@destinyishot what is the slope?

OpenStudy (anonymous):

What is the equation of the line, in point-slope form, that passes through the points (-7, -10) and (-6, -8)

OpenStudy (studystagram):

(9,1)-> x1,y1 & (4,16)->x2,y2 First find the slope which is by m=y2-y1/x2-x1 =16-1/4-9 m =-3 Equation of the point-slope given by y-y1=m(x-x1) y-1=-3(x-9) y-1=-3x+27 y=-3x+28

OpenStudy (studystagram):

(-7, -10) and (-6, -8) m=2 and equation is y=2x+4

OpenStudy (anonymous):

y + 10 = -2(x + 7) y - 10 = -2(x - 7) y + 10 = 2(x + 7) y - 10 = 2(x - 7) these are the choices i understand the m equals 2 but ....... so is it c or d

OpenStudy (anonymous):

d

OpenStudy (studystagram):

Hey, first understand the question and then try to learn it and I swear math is gonna be very easy but the option for the above quest is c :)

OpenStudy (studystagram):

Do you like want to understand the concept, I can help you with that :)

OpenStudy (anonymous):

its -8+10/-6+7 is 2/1 i understand but the second part i feel really blonde

OpenStudy (anonymous):

and yes i would i want to learn....

OpenStudy (triciaal):

The point-slope form of the equation of the line through (x1, y1) with slope m is y-y1 = m(x-x1)

OpenStudy (studystagram):

It's pretty easy see It's like you gotta use this formula-> y-y1=m(x-x1) and then sub in the values. y+10=2(x+7) y+10=2x+14 y=2x+4

OpenStudy (anonymous):

What is the equation of the line, in point-slope form, that passes through the points (-3, -1) and (-6, 8)? so the slope is -3 and im trying the other part

OpenStudy (anonymous):

y - 8 = -3(x + 6)

OpenStudy (anonymous):

im sorry i just dont understand where you got (x-x1)

OpenStudy (studystagram):

Listen first decide which points you're taking as x1,y1 and x2,y2. m is -3 yes. And the equation for the point-slope in y-y1=m(x-x1) The equation is that and then you gotta sub in the values like I've taken x1,y1 as -3,-1 so, y-1=-3(x+3) y=-3x-8

OpenStudy (anonymous):

y + 8 = -3(x - 6)

OpenStudy (studystagram):

Listen first, what did you take your x1,y1 & x2,y2 as?

OpenStudy (anonymous):

x1 is -3 y1 is -1 x2 is -6 y2 is 8

OpenStudy (triciaal):

@studystagram point-slope form is defined above do not rewrite in slope-intercept form @destinyishot there are 3 formats for writing the equation of a line

OpenStudy (anonymous):

What is the equation of the line, in point-slope form, that passes through the points (-3, -1) and (-6, 8)? y + 8 = -3(x - 6) y - 8 = -3(x + 6) y + 8 = 3(x - 6) y - 8 = 3(x + 6) ok heres the question and answers .......i get half of it i know is m equals -3 but now I'm so confused

OpenStudy (studystagram):

Okay you've taken it correct so you gotta sub in the values of x1,x2 and NOT X2,Y2. that's what you did there. You gotta do like this-> y-1=-3(x+3) so, opening up the bracket on R.H.S., y-1=-3x-9 y=-3x-8

OpenStudy (studystagram):

Wait, in your question they've taken (-6,8) as x1,y1 so yeah that's correct so the ans will be, y-8=-3(x+6) which is b

OpenStudy (studystagram):

You got it now? You should've given the options before lol so I would've taken the x1,y1 accordingly.

OpenStudy (anonymous):

i tryed ..... but yes thank you

OpenStudy (studystagram):

@triciaal, I really don't know what you meant by that, but I'm anyway answering by what I interpreted cuz you WILL get the ans in slope intercept form but you can change accordingly for ex- you can write also like y=-3x-8 OR -3x-y-8=0. Either way is correct and you WILL get the marks.

OpenStudy (anonymous):

What is the equation of the line, in point-slope form, that passes through the points (-1, 7) and (-2, 3)? y + 7 = -4(x - 1) y - 7 = -4(x + 1) y + 7 = 4(x - 1) y - 7 = 4(x + 1)

OpenStudy (studystagram):

@destinyishot Try and temme what you get the ans for the above question you've just posted and I will cross check with mine. If you don't understand, ask me.

OpenStudy (anonymous):

c

OpenStudy (studystagram):

It's d. Sign mistake. It's okay, you got the hang of it at least but you gotta be really attentive when there are -ve signs involved.

OpenStudy (anonymous):

ok i got it alright new question if your up for it.... What is the slope of the line that contains the points (-3, -2) and (4, -2)?

OpenStudy (anonymous):

well thanks anyway studystagram i really am grateful for your help............

Nnesha (nnesha):

it doesn't have the slope the slope is 0

OpenStudy (anonymous):

well would you use y2-y1/x2-x1 so it -4/7 right

Nnesha (nnesha):

yes i did but that is no -4

Nnesha (nnesha):

when you put y value -2-(-2) you should got 0

Nnesha (nnesha):

-2-(-2)/4-(-3) = -2+2/4+3 =0/7 m=0

OpenStudy (studystagram):

The slope is 0 as @Nnesha mentioned.

OpenStudy (anonymous):

ok did it back ward i think i understand

OpenStudy (anonymous):

- 4 undefined - 4/7 0 wait is it 0/7

OpenStudy (anonymous):

these are the answers

OpenStudy (studystagram):

It's 0, the ans.

OpenStudy (anonymous):

what about undefined

Nnesha (nnesha):

you got -4 bec you did -2-2 instead of -2-(-2)

OpenStudy (anonymous):

What is the slope of the line that contains the points (2, 6) and (-1, -6)? so 4

OpenStudy (studystagram):

Undefined is when you divide something by 0.

OpenStudy (studystagram):

Yeah the slope is 4. Well done hun! (:

Nnesha (nnesha):

yes good job @destinyishot

OpenStudy (triciaal):

@studystagram REPOST @studystagram point-slope form is defined above do not rewrite in slope-intercept form @destinyishot there are 3 formats for writing the equation of a line so you know 2 of the 3 the third is what the original question asked for any format will give the same line 1 point-slope 2 slope-intercept 3 standard

OpenStudy (anonymous):

What is the slope of the line that contains the points (-2, 8) and (-2, 1) so its 7/0 this is what you talking about undefined

OpenStudy (studystagram):

Yeah, that is undefined. Okay, I'm off to sleep. It's kinda late here :)

OpenStudy (anonymous):

ok thank you and from where......

OpenStudy (anonymous):

i hope your not some old creepy man in india

OpenStudy (studystagram):

Lol. I live in muscat and rest assured I'm a teen girl haha